# Quantum information with quantum-like bits

### v4

@@ III Quantum-like gates
To connect the usual definition of gates acting on an abstract Hilbert space to our QL formalism, it is convenient to introduce a map $U_{\mathrm{cb}}$ transforming the emergent eigenvectors onto the standard factorized computational basis. This map is defined by
$U_{\mathrm{cb}}=[V_{\mathrm{H}}\otimes\mathbb{1}_{N_{G}}]^{\otimes{N_{\mathrm{QL}}}},$ (III.1)
@@
$V_{\mathrm{H}}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (III.2)
denotes the standard Hadamard (H) gate. The action of $U_{\mathrm{cb}}$ becomes clear for ${N_{\mathrm{QL}}}=1$. Here,
$U_{\mathrm{cb}}\psi_{\uparrow}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{\downarrow}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix},$ (III.3)
where
$x_{M}=(x,\cdots,x)/\sqrt{M}\in\mathbb{C}^{M},\hskip 28.45274ptx\in\mathbb{C},\;M\in\mathbb{N}.$ (III.4)
We show in ?? how to construct the transformation $U_{\mathrm{cb}}$ in the more general case of non-regular graphs. The QL analog of any arbitrary quantum gate $V_{\mathrm{g}}$ is defined by the matrix transformation
$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (III.5)
Let us consider a few examples of ??. The QL maps of the identity operator and the three Pauli gates are, in the limit of exactly regular graphs,
$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.6a)
$U_{X}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.6b)
$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.6c)
$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.6d)
@@
As another example, the QL map of the Hadamard gate entangles the states of the factorized basis:
$U_{\mathrm{H}}\psi_{[\uparrow\downarrow]}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{[\uparrow\downarrow]}=\frac{1}{\sqrt{2}}(\psi_{\uparrow}\pm\psi_{\downarrow}).$ (III.8)
@@
$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}},$ (III.9)
transforming elements of the factorized basis as follows:
$U_{\mathrm{CN}}\psi_{\uparrow[\uparrow,\downarrow]}=\psi_{\uparrow[\uparrow,\downarrow]},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{\downarrow[\uparrow,\downarrow]}=\psi_{\downarrow[\downarrow,\uparrow]}.$ (III.10)
Note that, since a universal set of gates can be constructed by combining one-QL-bit gates and the CNOT, we can limit our analysis to this set of operations. Barenco et al. (1995)
@@
Up to this point, we have focused exclusively on the action of QL gates on states $\psi_{\bm{\sigma}}$ of the factorized basis. We can infer from those the transformation of any arbitrary state $\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\gamma_{\bm{\sigma}}\psi_{\bm{\sigma}}$, $\gamma_{\bm{\sigma}}\in\mathbb{C}$, thanks to linearity of the map ??. From an operational standpoint, it is however preferable to translate these state operations onto transformations applicable directed to the graph resources. This feature is key to translate our gate formalism into practical algorithms for classical computing resources and for experimental platforms of synchronizing electrical circuits. Ochs, Michaelis, and Roggendorf (2019) In this regard, we observe that the network transformation
$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{g}}=U_{\mathrm{g}}\mathcal{R}U_{\mathrm{g}}^{\dagger}$ (III.11)
transforms all eigenstates of a resource consistently to the gate $U_{\mathrm{g}}$ as defined in ??. This can be inferred from the spectral decomposition of ??, given by
$\mathcal{R}=\sum_{i=1}^{\tilde{N}}\lambda_{i}\psi_{i}\otimes\psi_{i}^{\dagger}.$ (III.12)
Here, $\tilde{N}=(2{N_{G}})^{N_{\mathrm{QL}}}$ denotes the total number of eigenstates of the resource $\mathcal{R}$, including both the emergent and the random states.
In the next ?? we shift the discussion towards the benefits provided by a suitable measure theory designed to add “quantumness” to classical synchronizing networks. This framework fulfills premises which can be key for a future quantification of QL advantage, as originally proposed in Ref. Khrennikov, 2019.
IV Quantum-like measurements
@@
In this section, we discuss criteria that allow to introduce a notion of QL interference of probabilities in our approach. This analysis highlights that genuine quantum mechanical systems are not strictly required to build QL stochastic models provided the notions of states, observables and measurement are defined according to a given set of prescriptions.
IV.1 Axiomatic definition
@@
$\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\alpha_{\bm{\sigma}}\psi_{\bm{\sigma}}.$ (IV.1)
@@
QL observables are represented by self-adjointed operators on $\mathcal{H}$. Their action on QL states, expressed by a linear map, defines a QL measurement.
@@
$\mathcal{F}$ $=\sum_{{\bm{\sigma},\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}}\left(\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}+\gamma^{*}_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right),\hskip 14.22636pt\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\in\mathbb{C},$ (IV.2)
@@
$P_{\psi}\left(\lambda_{\bm{\sigma}}^{\mathcal{F}}\right)=\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}=|\braket{f_{\bm{\sigma}},\psi}|^{2},$ (IV.3)
@@
$\Pi^{\mathcal{F}}_{\bm{\sigma}}=f_{\bm{\sigma}}\otimes f_{\bm{\sigma}}^{\dagger}=\left(\Pi^{\mathcal{F}}_{\bm{\sigma}}\right)^{2}.$ (IV.4)
@@
$\mathcal{F}f_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{F}}f_{\bm{\sigma}},\hskip 28.45274pt\mathcal{G}g_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{G}}g_{\bm{\sigma}}.$ (IV.5)
@@
$P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)=\frac{\left\lVert\Pi^{\mathcal{G}}_{\bm{\sigma}^{\prime}}\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}{\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}$ (IV.6)
@@
IV.2 Quantum probabilities in a QL model

### v5

@@ IV.1 Formalism
To connect the usual definition of gates acting on an abstract Hilbert space to our QL formalism, it is convenient to introduce a map $U_{\mathrm{cb}}$ transforming the emergent eigenvectors onto the standard computational basis. As that all elementary subrgaphs in our resources [????] are regular, this map can be expressed analytically by
$U_{\mathrm{cb}}=[V_{\mathrm{H}}\otimes\mathbb{1}_{N_{G}}]^{\otimes{N_{\mathrm{QL}}}},$ (IV.1)
@@
$V_{\mathrm{H}}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (IV.2)
denotes the matrix representation of the Hadamard (H) gate in the computational basis. The action of $U_{\mathrm{cb}}$ becomes clear for ${N_{\mathrm{QL}}}=1$, where
$U_{\mathrm{cb}}\psi_{\uparrow}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{\downarrow}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix},$ (IV.3)
and
$x_{M}=(x,\cdots,x)/\sqrt{M}\in\mathbb{C}^{M},\hskip 28.45274ptx\in\mathbb{C},\;M\in\mathbb{N}.$ (IV.4)
Note that, given that all the basis subgraphs are regular, the solution to ?? is independent of their local structure. We show in ?? that a solution for the map $U_{\mathrm{cb}}$ can be determined if the resources are not exactly regular, by solving numerically a linear algebra problem. We define the QL analog of any arbitrary quantum gate $V_{\mathrm{g}}$ by the matrix transformation
$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (IV.5)
To elucidate the action of ??, let us consider a few examples. The QL maps of the identity operator ($\mathcal{I}$) and the three Pauli gates (X,Y,Z) are
$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.6a)
$U_{X}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (IV.6b)
$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.6c)
$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (IV.6d)
@@
As another example, the QL map corresponding to the Hadamard gate transforms states in the factorized basis into superposition states:
$U_{\mathrm{H}}\psi_{\uparrow/\downarrow}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{\uparrow/\downarrow}=\frac{1}{\sqrt{2}}(\psi_{\uparrow}\pm\psi_{\downarrow}).$ (IV.8)
The formal approach proposed here applies equally to any arbitrary gates, including those outside the Clifford group. While gates outside the Clifford group break the stabilizer structure and require exponentially more resources compared to Clifford gates—which can be efficiently simulated classically—the distinction between these two classes of gates is not relevant to our framework. Our formalism is designed to be computationally exact, addressing the exponentially high cost associated with scaling QL resources, regardless of the type of gate involved. The action of the non-Clifford T-gate on the factorized basis is
$U_{\mathrm{T}}\psi_{\uparrow/\downarrow}=\frac{1}{2}\left[(1+\,\mathrm{e}^{i\pi/4})\psi_{\uparrow}+(1-\,\mathrm{e}^{i\pi/4})\psi_{\downarrow}\right].$ (IV.9)
@@
$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}}$ (IV.10)
transforms the elements of the factorized basis as follows:
$U_{\mathrm{CN}}\psi_{\uparrow[\uparrow,\downarrow]}=\psi_{\uparrow[\uparrow,\downarrow]},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{\downarrow[\uparrow,\downarrow]}=\psi_{\downarrow[\downarrow,\uparrow]}.$ (IV.11)
@@
Up to this point, we have focused exclusively on the action of QL gates on states $\psi_{\bm{\sigma}}$ of the factorized basis. We can infer from those the transformation of any arbitrary state $\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\gamma_{\bm{\sigma}}\psi_{\bm{\sigma}}$, $\gamma_{\bm{\sigma}}\in\mathbb{C}$, thanks to linearity of the map ??. From a practical standpoint, it is preferable to translate these state operations onto transformations applicable directed to the graph resources. This feature is a key requirement for a future experimental validation of our formalism with experimental resources. (An extended discussion on experimental resources is provided in the next ??). A straightforward approach in this regard is to define the unitary map
$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{g}}=U_{\mathrm{g}}\mathcal{R}U_{\mathrm{g}}^{\dagger}.$ (IV.12)
?? transforms all eigenstates of a resource consistently to the gate $U_{\mathrm{g}}$ as defined in ??. This property can be easily inferred by expanding $\mathcal{R}$ on the right-hand side of ?? via the spectral decomposition in ??. In the next ?? we provide concrete examples of ??, and we discuss how this network transformation affects the steady-state synchronizing solution of a system of oscillators.
IV.2 Connection to classical synchronization
Here, we explore the relationship between gate transformations of QL resources and the synchronization of classical many-body oscillator systems. Specifically, we focus on two examples of one-QL-bit gates. In these cases, the synchronization patterns of the classical system can be directly linked to the structure of the corresponding adjacency matrices. However, our approach is general and can be extended to arbitrary QL gates without any restrictions.
The Pauli-X (NOT) gate transforms an initial resource ?? as follows:
$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{X}}=\begin{pmatrix}A_{1}&-C\\ -C^{\top}&A_{2}\end{pmatrix}.$ (IV.13)
We show in panel a) of ??, the synchronization pattern in the dynamics of the Kuramoto model with correlation matrix $\mathcal{R}_{\mathrm{X}}$. In the transformation ??, the leading emergent eigenvector, associated to the largest eigenvalue $k+l$, is mapped from $\psi_{\uparrow}$ to $\psi_{\downarrow}$ [see ??]. The change of sign in the eigenvectors corresponds to a phase shift of $\Delta\phi=\pi$ between the first and the second half of the oscillators, as shown in the figure. In the case of the Hadamard gate, resources are transformed according to
$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{H}}=\begin{pmatrix}A_{1}+A_{2}+C+C^{\top}&A_{1}-A_{2}-C+C^{\top}\\ A_{1}-A_{2}+C-C^{\top}&A_{1}+A_{2}-C-C^{\top}\end{pmatrix}.$ (IV.14)
The sum and difference of the basis graphs $A_{1}$ and $A_{2}$ in ?? highlight the fundamental role of this gate in creating a superposition state. The structure of ?? encodes key information about the expected synchronization patterns in the dynamics of the corresponding Kuramoto model, as depicted in panel b) of ??. Dephasing emerges between the first and second halves of the oscillators, primarily driven by cancellation effects in the two off-diagonal blocks, where the term $A_{1}-A_{2}$ plays a dominant role. When $A_{1}\simeq A_{2}$ and $C\simeq C^{\top}$, the off-diagonal blocks approach zero, leading to completely uncorrelated phases between the two oscillator subgroups over extended timescales. Furthermore, the second half of the oscillators, labeled by indices $i\in\{{N_{G}},\cdots 2{N_{G}}\}$ exhibits weaker synchronization compared to the first one. This is due to the fact that the matrix norm of the lower diagonal block in ?? is, on average, smaller than that of the upper diagonal block. Consequently, the second group of oscillators exhibits weaker correlations compared to the first group.
In this section, we derived general principles for mapping network resources in a manner consistent with an arbitrary set of QL gates. Additionally, we presented numerical evidence supporting the connection between QL gates and the steady-state synchronization of classical oscillator resources. In the following ??, we explore specific examples of experimental resources that could be utilized to implement our formalism in practice.
V Experimental perspectives
So far, we have shown that the emergent states of synchronizing classical systems can be used as a basis to map the Hilbert space of a system of qubits. We discussed the primary role of emergent states in synchronization dynamics and the robustness of the latter in the presence of noise. In this section, we explore perspectives on implementing the proposed formalism using experimental physical resources. Additionally, we provide a high-level outline of computational and experimental validations that could support the further development of this framework.
For genuine quantum resources, the primary source of errors during information processing and measurement arises from uncontrolled noise. This noise can result from several factors, including external electromagnetic fields, unwanted interactions between qubits, fabrication defects, and imperfections in control protocols. Hou et al. (2016); Bengtsson et al. (2024); Funcke et al. (2022); Gupta and Nene (2021); Resch and Karpuzcu (2021) Addressing such errors typically requires multiple measurements to accurately resolve the final state of a quantum register after computations. This can introduce significant overhead in computing protocols and hinder the scalability of the approach to larger size. Error mitigation and decoding techniques, although conceptually scalable, can result in further limitations in size and performance. Bravyi et al. (2021); Ali et al. (2024) In our approach, error correction in the presence of noise translates to enhancing the robustness of long-time classical synchronization [see the discussion on panel b) of ??]. This ensures distinct and readily measurable transformations of the steady-state dynamics. Optimal synchronization can be achieved by numerically analyzing phase diagrams of synchronization indicators. For instance, in the case of the Kuramoto model, time-resolved order parameters offer clear insight into the synchronization strength. Basnarkov and Urumov (2007); Wang, Xu, and Zheng (2021) Parameter domains with the highest degree of synchronization can be identified as optimal candidates for experimental implementation.
In terms of resource complexity, the number of degrees of freedom required to model the synchronization dynamics of a system of ${N_{\mathrm{QL}}}$ QL bits scales exponentially with system size. Although this scaling does not impose fundamental constraints on computational studies or experimental implementations of the approach, it does limit the practical feasibility of large-scale QL algorithms. Here, we discuss a few experimental platforms that are suitable candidates for implementing QL information processing within these scaling constraints. Additionally, we identify the classes of algorithms that can be most effectively analyzed by those resources.
Implementing coupled LC circuits is a straightforward approach to implement macroscopic experimental resources that exhibit classical synchronization. Allam et al. (2006); Barbi (2021) Numerical analyses from Ref. Maffezzoni et al., 2015 show that this class of systems can be effectively approximated by Kuramoto-type models for medium-scale systems of 60 oscillators. The couplings between the model oscillators are tuned by adjusting the transconductance values of the circuit transistors.
Spin-torque oscillators (STO) emerge as another promising candidate for implementing QL information processing. Raychowdhury et al. (2019) STOs are nonlinear magnetic oscillators that can be integrated into nanoscale devices operating at microwave frequencies. They have the ability to synchronize in both frequency and phase with external oscillatory signals as well as with other STOs. The full magnetization dynamics of these systems, defined by the Landau-Lifshitz-Gilbert-Slonzewski (LLGS) equation, can be simplified to the Kuramoto model for small-amplitude oscillations. Belanovsky et al. (2013); Kanao et al. (2019) Ref. Flovik, Macià, and Wahlström, 2016 demonstrates the experimental feasibility of modeling synchronizing classical systems with up to a few hundred oscillators. Beyond this scale, technical constraints arising from finite-size effects restrict the extent of achievable synchronization.
Metal-to-insulator transition (MIT) devices leverage materials that can switch between insulating and metallic (conductive) states in response to external conditions such as temperature, pressure, or electrical stimuli. Kim et al. (2010) Research on small-scale MIT circuits has demonstrated their fundamental ability to produce classical synchronization behavior through voltage-triggered oscillatory dynamics. Kim et al. (2010); Parihar et al. (2015); Raychowdhury et al. (2019) One limitation of current MIT devices is their low oscillation frequency, typically below 1 MHz, which hinders the ability to achieve fine-resolution synchronization and detect subtle variations in coupled oscillator behavior. However, ongoing advancements in scaling down these devices are expected to increase their oscillation frequencies, enabling more precise and time-resolved synchronization measurements. Cotter et al. (2014)
Building on the experimental perspectives outlined above, it appears feasible to validate our QL-bit theory using networks of up to a hundred oscillators. Such an experimental setup would enable the implementation of proof-of-concept algorithms on systems with ${N_{\mathrm{QL}}}=2$ or ${N_{\mathrm{QL}}}=3$ QL-bits, where each QL-bit comprises a small network of 555510101010 oscillators. A particularly intriguing class of quantum information processes for systems of this scale includes those that generate entanglement, such as the formation of Bell states (for ${N_{\mathrm{QL}}}=2$) and their higher-order generalizations, the Greenberger–Horne–Zeilinger (GHZ) states. Greenberger, Horne, and Zeilinger (1989); Erhard, Krenn, and Zeilinger (2020) This analysis has the potential to provide valuable insights into fundamental questions surrounding the classical-to-quantum correspondence. In particular, it could help clarify the role of non-emergent states in classical synchronizing systems as potential hidden variables within a QL-based framework for analyzing entanglement. Scholes (2025b)
We conclude this section by presenting a practical perspective on implementing an experimental scheme to test our approach:
  1. 1.
    Engineer a classical synchronizing system: Design a system with two-body correlations represented by a graph, ensuring that its leading emergent state corresponds to the desired initial state of a QL algorithm. For example, the factorized state $\psi_{\uparrow}$ can be associated with a Cartesian-product network structured as described in ??.
  2. 2.
    Initialize oscillator phases: Assign arbitrary initial conditions to the oscillator phases, ensuring they adhere to the constraint that the system’s long-time steady state achieves the expected synchronization.
  3. 3.
    Allow the classical dynamics to equilibrate: Let the system evolve until it reaches a steady state. This step requires time-resolved measurements to ensure the system has sufficient time to reach a long-time synchronized state.
  4. 4.
    Perform gate operations: Each gate operation consists of two phases: (a) transforming the correlations between the degrees of freedom and (b) allowing the system to synchronize and relax to a new emergent state.
  5. 5.
    Measure the final state: Apply projection operators to extract the state of the final resource after all gate operations have been implemented.
The proposed scheme can be tested numerically by measuring multi-time correlation functions. This process involves a series of time evolution operators defined by gate-transformed resources at different stages of a given QL algorithm. Preliminary analytical work on the Kuramoto dynamics is discussed in ?? [see ??].
Figure 6: a): Time evolution of a Kuramoto model where the adjacency matrix is transformed according to the mapping in ?? for the Pauli-X gate (g = X). b): Similar to panel (a), but for the Hadamard gate (g = H). The same model parameters as defined in ?? are used here.
VI Quantum-like interfence
@@
In this final section, we discuss criteria that allow to introduce a notion of QL interference of probabilities in our approach. This analysis highlights that genuine quantum mechanical systems are not strictly required to build QL stochastic models provided the notions of states, observables and measurement are defined according to a given set of prescriptions.
VI.1 Axiomatic definition
@@
$\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\alpha_{\bm{\sigma}}\psi_{\bm{\sigma}}.$ (VI.1)
@@
QL observables are represented by self-adjoint operators on $\mathcal{H}$. Their action on QL states, expressed by a linear map, defines a QL measurement.
@@
$\mathcal{F}$ $=\sum_{{\bm{\sigma},\bm{\sigma}^{\prime}\in\bm{\mathcal{S}}}}\left(\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}+\gamma^{*}_{\bm{\sigma},\bm{\sigma}^{\prime}}\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right),\hskip 14.22636pt\gamma_{\bm{\sigma},\bm{\sigma}^{\prime}}\in\mathbb{C},$ (VI.2)
@@
$P_{\psi}\left(\lambda_{\bm{\sigma}}^{\mathcal{F}}\right)=\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}=|\braket{f_{\bm{\sigma}},\psi}|^{2},$ (VI.3)
@@
$\Pi^{\mathcal{F}}_{\bm{\sigma}}=f_{\bm{\sigma}}\otimes f_{\bm{\sigma}}^{\dagger}=\left(\Pi^{\mathcal{F}}_{\bm{\sigma}}\right)^{2}.$ (VI.4)
@@
$\mathcal{F}f_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{F}}f_{\bm{\sigma}},\hskip 28.45274pt\mathcal{G}g_{\bm{\sigma}}=\lambda_{\bm{\sigma}}^{\mathcal{G}}g_{\bm{\sigma}}.$ (VI.5)
@@
$P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)=\frac{\left\lVert\Pi^{\mathcal{G}}_{\bm{\sigma}^{\prime}}\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}{\left\lVert\Pi^{\mathcal{F}}_{\bm{\sigma}}\psi\right\rVert^{2}}$ (VI.6)
@@
VI.2 Quantum probabilities in a QL model